GDP in the United States was $16,150 billion in 2012, and $19,400 billion in 2017. Over that same five years, population rose from 314 million people to 326 million, and prices rose by a total of 7.8%. The average annual rate of real per-capita growth from 2012-2017 was:

a) 1.0%

b) 1.4%

c) 1.5%

d) 2.1%

e) 7.3%
Answer #1

Answer

the correct answer is (b) 1.4%

Let 2012 be a base year and hence price in 2012 = 100 and as it grows by total of 7.8%. Thus price level in 2017 = 100 + (7.8/100)*100 = 107.8

Real GDP = (nominal GDP/Price level)*100

=> Real GDP per capita = ((nominal GDP/Price level)*100)/Population.

Hence, Real GDP per capita in 2012 = ((16150 billion/100)*100)/314 million = 51433.12

Real GDP per capita in 2017 = ((19400 billion/107.8)*100)/326 million = 55203.34

Average annual growth rate = (New value / Initial Value)1/n - 1

where n = time in years = 5

=> Average annual growth rate of real GDP per capita = (55203.34 / 51433.12)1/5 - 1 = 0.0142 ~ 1.4%(approx.).

Hence, the correct answer is (b) 1.4%